Find the escape speed for a rocket leaving Titan. The acceleration of gravity on Titan is 0.130 times that on earth, and Titans radius is 0.407 time the radius of the earth X m/s
Answer #1

Here ,

for the escape velocity

escape velocity = sqrt(2 * GM/R)

escape velocity = sqrt(2 * GM/R^2 * R)

escape velocity = sqrt(2 * g * R)

escape velocity = sqrt(2 * 0.130 * 9.8* 0.407 * 6371 *10^3)

escape velocity = 2570 m/s

the escape velocity on the planet is 2570 m/s